OpenStudy (anonymous):

Need help on how to solve this problem Determine whether or not E and F are independent or dependent. P(E) = 0.4, P(F) = 0.5, P(E ∩ F) = 0.2 a.) independent b.)dependent

myininaya (myininaya):

Well does P(E)*P(F)=P(E intersect F)?

OpenStudy (anonymous):

If P(E ∩ F) = P(E) x P(F) , then they are independent

OpenStudy (anonymous):

Alright, Thank You